When a Dc current goes through a coil, it will create a magnetic flux, after a while this flux will be stable and the resistance goes to zero because of a short-circuit and the wire is heated.

This was the explanation for what happens when a dc current is connected to a coil. My question is, when the current start flowing in the wire heat is going to be released and the resistance will increase.. Is this before the short-circuit to occur? Because after short-circuit the resistance is zero but before that the resistance increase and the wire get heated. right?

Why does exactly the short-circuit happens when the magnetic flow is stable?
Answer #1

For a coil V=L di/dt is the voltage equation.since there is no change of current in DC ,di/dt=0 except during the transient period.Then V=0.This means short circuit.So high currents will flow and burn out the winding. There will be no induced currents in the circuit as magnetic flux is not changing through coil. It can also be explained in another way.

for a coil Z=R+jωL. ω= 2*pi*f,For DC f=0,so ω=0 ,Z=R.For a coil L will be more and R will be less.So Z will be very less and very high currents will flow and damage the circuit.

I hope this clears your doubt.

Thanks !